OpenStudy (anonymous):

Which point lies in the solution set for the following system of inequalities? y > x + 4 y > -2x + 2 (-4, 0) (-3, 2) (2, -1) (0, 6)

OpenStudy (anonymous):

this is when you start subbing your points into the inequality to see if they are true. For it to be the solution, your points have to be true in both inequalities. I will do the first one for you... (-4,0) x = -4, y = 0 y > x + 4 0 > -4 + 4 0 > 0 (not true because 0 = 0) There is no need to try the other inequality if the first one is not true. you want to try your second set of points ?

OpenStudy (anonymous):

Yeah, I'll try.

OpenStudy (anonymous):

is it "a"? @kelliegirl33

OpenStudy (anonymous):

no....I did the first one and it is not the solution

OpenStudy (anonymous):

it's not true right? 0>0 is invalid right?

OpenStudy (anonymous):

lets try the 2nd one... (-3,2) y > x + 4 2 > -3 + 4 2 > 1 (true) y > -2x + 2 2 > -2(-3) + 2 2 > 6 + 2 2 > 8 (not true) 3rd set of points... (2,-1) y > x + 4 -1 > 2 + 4 -1 > 6 (no true) 4th set of points... (0,6) y > x + 4 6 > 0 + 4 6 > 4 (true) y > -2x + 2 6 > -2(0) + 2 6 > 0 + 2 6 > 2 (true) the 4th one is your answer. Do you understand ?

OpenStudy (anonymous):

if you have any questions, please ask

OpenStudy (anonymous):

Thank you so much! Can you show me how you did the equation step by step because I suck at math!

OpenStudy (anonymous):

sure... your points are (0,6) with x being 0 and y being 6 now we take our inequalities : y > x + 2 (sub 0 in for x and 6 in for y) 6 > 0 + 2 6 > 2 (this is correct because 6 is greater then 2) > means greater and < means less ( I learned it by saying the alligator eats the biggest number.....12 > 6....the alligators mouth is open towards the biggest number) second inequality.... y > -2x + 2 sub in 0 for x and 6 for y 6 > -2(0) + 2 6 > 0 + 2 6 > 2 (true) -- alligator eats the biggest number, which is 6...lol

OpenStudy (anonymous):

did that help ?

OpenStudy (anonymous):

Thank you so much! you're so nice for helping me, appreciate it.

OpenStudy (anonymous):

anytime :)

OpenStudy (anonymous):

can you help me with some more?

OpenStudy (anonymous):

what you got ?

OpenStudy (anonymous):

For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? -x + 2y = -6 3x + y = 8 -2y + 6 2y - 6 2y + 6 -2y - 6